# Schrödinger Symmetry in Spherically-symmetric Static Mini-superspaces with Matter Fields

### v1

@@ Abstract
Schrödinger symmetry emerged in a “fluid limit” from a full superspace to several mini-superspace models. We consider two spherically-symmetric static mini-superspace models with matter fields and verify the robustness of this emergent symmetry at the classical level: (i) Maxwell field with cosmological constant and (ii) $n$ massless scalar fields. We develop a method based on canonical transformations and show that: for model (i), 3D Schrödinger symmetry emerges, and the solution is the (anti-)de Sitter Reissner-Nordström spacetime, and for model (ii), (2+n)(2+n)D Schrödinger symmetry appears, and the solution is a generalized Janis–Newman–Winicour spacetime and its “interior”, a Kantowski-Sachs type closed universe. In the matter decoupling limit, both cases lead to 2D Schrödinger symmetry in different lapse functions and mini-superspace coordinates, which implies the covariance of Schrödinger symmetry. Finally, we propose a physical interpretation of the symmetry under Hamiltonian constraints $H$ and explain it with examples: Symmetry generators commuting with $H$ map a solution to another one, while those non-commuting with $H$ generate a new theory with the Schrödinger symmetry and the transformed configuration is a solution to the new theory. These support the robustness of the emergence of Schrödinger symmetry and open new possibilities for exploring quantum dynamics of matter and gravity based on the symmetry.
@@ I Introduction
Symmetry has played a significant role in physics. In the quest for a theory of quantum gravity, where the fundamental principles are still unknown, symmetry should serve as a guiding principle. The characteristic symmetry of gravity is general covariance, the gauge symmetry of spacetime, which leads to the background independence and dynamics of gravity. Here, the Hamiltonian constraint $\mathcal{H}=0$ is the master key; it generates diffeomorphism Dirac , the dynamics of gravity ultimately reduce to it Peres1962 ; DeWitt:1967yk ; osti_4124259 ; Gerlach:1969ph ; Moncrief:1972cx , and it plays a central role in various crucial problems, such as the so-called “time problem” Kuchar:1991qf ; Isham:1992ms ; Hoehn:2020epv and the quantization of gravity DeWitt:1967yk ; osti_4124259 . Thus, the Hamiltonian constraint should be the essential ingredient in exploring the symmetry of (quantum) gravity.
@@
Model I describes (anti-)de Sitter Reissner-Nordström (AdS-RN) solutions, which reduce to the Schwarzschild solutions in the limit of vanishing electric charge and cosmological constant. It thus serves as a natural first step toward incorporating matter into black hole mini-superspace models. Model II leads to the Janis–Newman–Winicour (JNW) solution Janis:1968zz , which asymptotes to the Schwarzschild geometry at $r\to\infty$ but lacks an event horizon and instead possesses a naked singularity, and it also provides a basic setup for relational formalism using a scalar field as a “clock” Ashtekar:2011ni ; Domagala:2010bm .
In a spherically symmetric spacetime, the areal radius $r$ is well-defined, and it should be natural to consider foliations of spacetime in terms of $r$ or an associated parameter $\tau$. Then, integrating over time (due to the static nature) and over spherical directions (due to spherical symmetry) leaves only $\tau$. This parameter serves as “time”, and we obtain the Hamiltonian $H=N\mathcal{H}$. Here, $N$ is the lapse function, and $\mathcal{H}$ is the Hamiltonian constraint that guarantees the reparametrization invariance for $\tau$. Thus, the two mini-superspace models are a kind of one-dimensional mechanical system with $H=0$ (“black-hole mechanics” Cavaglia:1994yc ; Bodendorfer:2019cyv ; Achour:2021dtj ; BenAchour:2022fif ; BenAchour:2023dgj ; Perez:2023jrq ; Livine:2025soz ). We study the symmetry of such systems.
In Sec.II, we see that for a free-particle system, Schrödinger symmetry can be obtained easily by calculating the algebra of the Noether charges. Here, symmetry means that a Hamiltonian does not change by the transformation (while a more physical meaning is discussed below). Our strategy for obtaining the symmetry in the mini-superspace models is to choose the lapse function $N$ properly and perform canonical transformations so that the Hamiltonian $H$ takes the same form as that of the free particle system. Then, the Schrödinger algebra can be constructed (without imposing $H=0$), and the Schrödinger symmetry appears in terms of the new canonical variables. Here, the solution configuration is represented as a uniform linear motion in the mini-superspace with the momentum constrained by $H=0$, which can be translated into the spacetime geometry through the canonical transformation.
@@
In Sec.IV, we analyze Model I, where the presence of the Maxwell field $A_{0}(r)$ increases the dimension of the mini-superspace by one. By applying our strategy, we show that the 3D Schrödinger symmetry appears. The 2D Schrödinger symmetry is embedded as a subalgebra, and in the limit of vanishing the electric charge, the 3D symmetry reduces to the 2D one. (A)dS-RN metrics can be obtained as the solution in a consistent manner.
@@
In Sec.VI, we discuss the physical meaning of the obtained symmetry under the Hamiltonian constraint $H=0$. First, the generators $G$ of the Schrödinger symmetry obtained for a certain choice of $N$ satisfy $\{G,H\}+\partial_{\tau}G=0$. Momenta $p_{i}$ and angular momenta $j_{ij}$ commute with $H$, and we can show that such generators transform a solution to another one satisfying $H=0$ (dynamical symmetry). As examples, we construct symmetry transformations that map the RN solution to the Schwarzschild one and those that map the JNW solution to the Schwarzschild one. On the other hand, Galilean boost $b^{i}$, dilatation $q_{0}$, and special conformal transformation $q_{-}$ do not commute with $H$, which prevents interpretation as a dynamical symmetry consistent with $H=0$. So, we suggest one possible interpretation: such a non-commutative symmetry transforms the theory itself, $H\to H_{\rm New}$, and the transformed configuration is a physical solution of the new theory with $H_{\rm New}=0$. We will demonstrate this concretely: in Model I, as the effect of adding a two-dimensional cosmological term; in Model II, as the effect of adding a new scalar field. This provides one interpretation of Noether’s theorem in a theory with the Hamiltonian constraint $H\approx 0$ and symmetry generators $G$ that do not commute with $H$.
@@ II.1 Free particle
which is the condition for the Hamiltonian to be invariant. [^2] If no Hamiltonian constraint exists, this leads to dynamical symmetries. See Sec.VI for more details. That is, a trajectory generated by such a function $G$ from a solution $(x^{i}(t),p_{i}(t))$,
@@
also satisfies the canonical equation as another solution with different values of integration constants, where $\epsilon$ is a small parameter. On the other hand, such a generator $G$ is conserved in on-shell:
@@
Generally, in the case of the free-particle Hamiltonian, symmetry generators $G$ transform a uniform linear motion with certain values of $(\alpha_{i},\beta^{i})$ to another one with different ones $(\alpha^{\prime}_{i},\beta^{\prime i})$. (We will provide a more explicit argument in Sec.VI.) In particular, the momentum $p_{i}$ and the Galilean boost generator $b^{i}$ defined by
@@
Generically, gravitational (mini-)superspaces have a Lorentizan signature DeWitt:1967yk . Indeed, a Lorentzian version of the Schrödinger algebra will appear in our models below.
@@ II.2 Canonical-transformation Method
An interesting and important question for understanding the robustness of the emergent symmetry of gravitational systems is: under which condition can the Schrödinger symmetry appear in general? Although finding a complete answer would be non-trivial, there is one related result based on Eisenhart–Duval (ED) lift method; one considers a “lifted” configuration space endowed with a lifted metric and can show that when the lifted space is conformally flat, the Schrödinger symmetry is obtained cccc1220-d9a2-3162-a3c5-a8782bbc8c8b ; universe9120503. Instead of using this method, we will consider more general (but not completely general yet) cases in which the ED method may not be useful, and find a sufficient condition for the Schrödinger symmetry to appear.
@@
where $\tilde{g}^{ij}$ is a constant metric. Note here that canonical transformations preserve the symplectic structure: $\{x^{i},p_{j}\}=\delta^{i}_{j}\to\{\tilde{x}^{i},\tilde{p}_{j}\}=\delta^{i}_{j}$. Therefore, we can apply the same procedure as in the previous subsection to construct the generators of the Schrödinger algebra. Formally, they are given by replacing $(x^{i},p_{i})$ with $(\tilde{x}^{i},\tilde{p}_{i})$ in the generators (9) and (11):
@@ III.1 X-coordinate in the mini-superspace
$S_{g}[g_{\mu\nu}]=\frac{1}{16\pi G\hbar}\int_{\mathcal{M}}d^{4}x\sqrt{-g}(R-2\Lambda),$ (27)
@@
for functions on the phase space: $f_{1}(X,P),~f_{2}(X,P)$. We then obtain the Hamiltonian:
@@
$\mathcal{H}_{g}$ is the Hamiltonian constraint, which must be set to zero in order to gain the physical solution: ${\cal H}_{g}\approx 0$. Here, $\approx$ is the weak equality that holds after calculating Poisson brackets, as usual.
We now fix the gauge as
@@
where $a,b=+,$-, and the mini-superspace is “Lorentzian”, $\rm{det}\mathcal{G}_{ab}<0$, as consistent with the general result DeWitt:1967yk . Note here that the lifted space is not conformally flat due to $\Lambda\neq 0$, which does not allow one to apply the ED method and directly find Schrodinger symmetries universe9120503.
@@
According to the explanation in the previous subsection, our first step is to find conserved quantities $f_{a}(X,P)$ and identify them with new momenta $\widetilde{P}_{a}(X,P)$. Noting that $\{H_{g},H_{g}\}=\{H_{g},P_{+}\}=0$, we can write (35) as
@@
$\widetilde{X}^{+}=X^{+}-\frac{\epsilon c^{2}l_{p}^{2}}{12l_{\Lambda}^{2}}\frac{(X^{-})^{3}}{(P_{+})^{2}},~~~\widetilde{X}^{-}=X^{-}$ (41)
@@
Thus, we complete the canonical transformation, and the Hamiltonian (35) becomes, in the gauge (34), the desired form:
@@
We conclude this derivation with three remarks. First, the above symmetry holds for the specific choice of the the lapse function (34). Second, the Hamiltonian constraint is not used to derive the symmetry. Third, the constraint itself appears as one of the generators, $\widetilde{Q}_{+}=H_{g}+\frac{cl_{p}}{2l_{s}^{2}}$. These three are intertwined in understanding the physical meaning of the symmetry (see Sec.VI).
@@
Before closing this section, we here show that the solution of the Hamiltonian (35) is uniquely given by the (anti)de Sitter Schwarzschild metric, which provides a physical aspect of this system. In the derivation, we impose the Hamiltonian constraint $H_{g}\approx 0$ as late as possible to make its role clear.
@@
Combining these and the gauge condition (34), we can calculate
@@
Here, the formula (53) plays the role of the energy of the system. Indeed, in the case of $\Lambda=0$ and $\epsilon=-1$, it agrees with the Komar mass Wald:1984rg (see Appendix A). Note that $M$ commutes with $H_{g}$ (43) and thus is gauge invariant.
@@ IV Maxwell Field
Here, $\mathcal{H}_{gE}$ is the Hamiltonian constraint, which becomes $\mathcal{H}_{gE}\approx 0$ on shell, and the canonical variables $(X^{+},X^{-},X^{A};P_{+},P_{-},P_{A})$ satisfy the relation (31).
@@
which indeed agrees with the standard definition of electric charge (see Appendix A). In the last line, we employed the Hamiltonian constraint (from (69)):
@@ V Multiple Massless Scalar Fields
As another case with matter fields, we discuss the mini-superspace with $n$ massless scalar fields, whose solution corresponds to a generalized Janis–Newman–Winicour (JNW) solution Janis:1968zz and its “interior” Chiou:2008eg , and we show that $(2+n)$ D-Schrödinger symmetry appears. An interesting point here is that the symmetry is obtained by using a different coordinate system $Y^{a}$ in the mini-superspace than the previous one $X^{a}$, together with a different choice of gauge $N$.

### v2

@@ Abstract
Schrödinger symmetry emerged in a “fluid limit” from a full superspace to several mini-superspace models. We consider two spherically-symmetric static mini-superspace models with matter fields and verify the robustness of this emergent symmetry at the classical level: (i) Maxwell field with cosmological constant and (ii) $n$ massless scalar fields. We develop a method based on canonical transformations and show that: for model (i), 3D Schrödinger symmetry emerges, and the solution is the (anti-)de Sitter Reissner-Nordström spacetime, and for model (ii), (2+n)(2+n)D Schrödinger symmetry appears, and the solution is a generalized Janis-Newman-Winicour spacetime and its “interior”, a Kantowski-Sachs type closed universe. In the matter decoupling limit, both cases lead to 2D Schrödinger symmetry in different lapse functions and mini-superspace coordinates, which implies the covariance of Schrödinger symmetry. Finally, we propose a physical interpretation of the symmetry under Hamiltonian constraints $H$ and explain it with examples: Symmetry generators commuting with $H$ map a solution to another one, while those non-commuting with $H$ generate a new theory with the Schrödinger symmetry and the transformed configuration is a solution to the new theory. These support the robustness of the emergence of Schrödinger symmetry and open new possibilities for exploring quantum dynamics of matter and gravity based on the symmetry.
@@ I Introduction
Symmetry has played a significant role in physics. In the quest for a theory of quantum gravity, where the fundamental principles are still unknown, symmetry should serve as a guiding principle. The characteristic symmetry of gravity is general covariance, the gauge symmetry of spacetime, which leads to the background independence and dynamics of gravity. Here, the Hamiltonian constraint $\mathcal{H}=0$ is the master key; it generates diffeomorphism Dirac , the dynamics of gravity ultimately reduce to it Peres1962 ; DeWitt:1967yk ; osti_4124259 ; Gerlach:1969ph ; Moncrief:1972cx , and it plays a central role in various crucial problems, such as the so-called “problem of time” Kuchar:1991qf ; Isham:1992ms ; Hoehn:2020epv and the quantization of gravity DeWitt:1967yk ; osti_4124259 . Thus, the Hamiltonian constraint should be the essential ingredient in exploring the symmetry of (quantum) gravity.
@@
Model I describes (anti-)de Sitter Reissner-Nordström (AdS-RN) solutions, which reduce to the Schwarzschild solutions in the limit of vanishing electric charge and cosmological constant. It thus serves as a natural first step toward incorporating matter into black hole mini-superspace models. Model II leads to the Janis-Newman-Winicour (JNW) solution Janis:1968zz , which asymptotes to the Schwarzschild geometry at $r\to\infty$ but lacks an event horizon and instead possesses a naked singularity, and it also provides a basic setup for relational formalism using a scalar field as a “clock” Blyth:1975is ; Halliwell:1984eu ; Ashtekar:2011ni ; Domagala:2010bm .
In a spherically symmetric spacetime, the areal radius $r$ is well-defined, and it should be natural to consider foliations of spacetime in terms of $r$ or an associated parameter $\tau$. Then, integrating over time (due to the static nature) and over spherical directions (due to spherical symmetry) leaves only $\tau$. This parameter serves as “time”, and we obtain the Hamiltonian $H=N\mathcal{H}$. Here, $N$ is the lapse function, and $\mathcal{H}$ is the Hamiltonian constraint that guarantees the reparametrization invariance for $\tau$. Thus, the two mini-superspace models are a kind of one-dimensional mechanical system with Hamiltonian constraints $H=0$ (“black-hole mechanics” Cavaglia:1994yc ; Bodendorfer:2019cyv ; Achour:2021dtj ; BenAchour:2022fif ; BenAchour:2023dgj ; Perez:2023jrq ; Livine:2025soz ). We study the symmetry of such systems.
In Sec.II, we see that for a free-particle system, Schrödinger symmetry can be obtained easily by calculating the algebra of the Noether charges. Here, symmetry means that a Hamiltonian does not change by the transformation (see below for a more physical meaning). Our strategy for obtaining the symmetry in the mini-superspace models is to choose the lapse function $N$ properly and perform canonical transformations so that the Hamiltonian $H$ takes the same form as that of the free particle system. Then, the Schrödinger algebra can be constructed (without imposing $H=0$), and the Schrödinger symmetry appears in terms of the new canonical variables. Here, the solution configuration is represented as a uniform linear motion in the mini-superspace with the momentum constrained by $H=0$, which can be translated into the spacetime geometry through the canonical transformation.
@@
In Sec.IV, we analyze Model I, where the presence of the Maxwell field $A_{0}(r)$ increases the dimension of the mini-superspace by one. By applying our strategy, we show that the 3D Schrödinger symmetry appears. The 2D Schrödinger symmetry is embedded as a subalgebra, and in the limit of vanishing the electric charge, the 3D symmetry reduces to the 2D one. (A)dS-RN metrics can be obtained as the solution in a consistent manner. In Appendix A, the formulae of mass and electric charge are discussed.
@@
In studying these models, we find one interesting result. In both models, the ADM energy $M$, represented in different canonical coordinates, commutes with the Hamiltonian constraints and thus is gauge invariant. We can construct a gauge-invariant observable $T$ that is associated with the time coordinate at infinity and canonically conjugate to $M$, $\{T,M\}=1$.
In Sec.VI, we discuss the physical meaning of the obtained symmetry under the Hamiltonian constraint $H=0$. First, the generators $G$ of the Schrödinger symmetry obtained for a certain choice of $N$ satisfy $\{G,H\}+\partial_{\tau}G=0$. Momenta $p_{i}$ and angular momenta $j_{ij}$ commute with $H$, and we can show that such generators transform a solution to another one satisfying $H=0$ (dynamical symmetry). As examples, we construct symmetry transformations that map the RN solution to the Schwarzschild one and those that map the JNW solution to the Schwarzschild one. On the other hand, Galilean boost $b^{i}$, dilatation $q_{0}$, and special conformal transformation $q_{-}$ do not commute with $H$, which prevents interpretation as a dynamical symmetry consistent with $H=0$. So, we suggest one possible interpretation: such a non-commutative symmetry transforms the theory itself, $H\to H_{\rm New}$, and the transformed configuration is a physical solution of the new theory with $H_{\rm New}=0$. We will demonstrate this concretely: in Model I, as the effect of adding a two-dimensional cosmological term, and in Model II, as the effect of adding a new scalar field. This provides one interpretation of Noether’s theorem in a theory with the Hamiltonian constraint $H\approx 0$ and symmetry generators $G$ that do not commute with $H$. In Appendix B, the symmetry is studied in the action principle.
@@ II.1 Free particle
which is the condition for the Hamiltonian to be invariant. That is, a trajectory generated by such a function $G$ from a solution $(x^{i}(t),p_{i}(t))$,
@@
also satisfies the canonical equation as another solution with different values of integration constants, where $\epsilon$ is a small parameter. [^2] This is a dynamical symmetry if no Hamiltonian constraint exists. See Sec.VI for more details. On the other hand, such a generator $G$ is conserved in on-shell:
@@
In the case of the free-particle Hamiltonian (2), symmetry generators $G$ transform a uniform linear motion (5) with certain values of $(\alpha_{i},\beta^{i})$ to another one with different ones $(\alpha^{\prime}_{i},\beta^{\prime i})$. (We will provide a more explicit argument in Sec.VI.) In particular, the momentum $p_{i}$ and the Galilean boost generator $b^{i}$ defined by
@@
Generically, gravitational (mini-)superspaces have a Lorentizan signature DeWitt:1967yk . Indeed, a Lorentzian version of Schrödinger algebra will appear in our models below.
@@ II.2 Canonical-transformation Method
An interesting and important question for understanding the robustness of the emergent symmetry of gravitational systems is: under which condition can Schrödinger symmetry appear in general? Although finding a complete answer would be non-trivial, there is one related result based on Eisenhart–Duval (ED) lift method; one considers a “lifted” configuration space endowed with a lifted metric and can show that when the lifted space is conformally flat, the Schrödinger symmetry is obtained cccc1220-d9a2-3162-a3c5-a8782bbc8c8b ; BenAchour:2022fif . Instead of using this method, we will consider more general (but not completely general yet) cases in which the ED method may not be useful, and find a sufficient condition for the Schrödinger symmetry to appear.
@@
where $\tilde{g}^{ij}$ is a constant metric. Note here that canonical transformations preserve the symplectic structure: $\{x^{i},p_{j}\}=\delta^{i}_{j}\to\{\tilde{x}^{i},\tilde{p}_{j}\}=\delta^{i}_{j}$. Therefore, we can apply the same procedure as in the previous subsection to construct the generators of Schrödinger algebra. Formally, they are given by replacing $(x^{i},p_{i})$ with $(\tilde{x}^{i},\tilde{p}_{i})$ in the generators (9) and (11):
@@ III.1 X-coordinate in the mini-superspace
$S_{g}[g_{\mu\nu}]=\frac{1}{16\pi G\hbar}\int_{\mathcal{M}}d^{4}x\sqrt{-g}(R-2\Lambda).$ (27)
@@
for functions $f_{1},f_{2}$ on the phase space. We then obtain the Hamiltonian:
@@
$\mathcal{H}_{g}$ is the Hamiltonian constraint, which must be set to zero in order to gain the physical solution: ${\cal H}_{g}\approx 0$. Here, $\approx$ is the weak equality that holds after calculating Poisson brackets, as usual Dirac .
We now fix the lapse as
@@
where $a,b=+,$-, and the mini-superspace is “Lorentzian”, $\rm{det}\mathcal{G}_{ab}<0$, as consistent with the general result DeWitt:1967yk . Note here that the lifted space is not conformally flat due to $\Lambda\neq 0$, which does not allow one to apply the ED method and directly find Schrodinger symmetries BenAchour:2022fif .
@@
According to the explanation in the previous section, our first step is to find conserved quantities $f_{a}(X,P)$ and identify them with new momenta $\widetilde{P}_{a}(X,P)$. Noting that $\{H_{g},H_{g}\}=\{H_{g},P_{+}\}=0$, we can write (35) as
@@
$\widetilde{X}^{+}=X^{+}-\frac{\epsilon c^{2}l_{p}^{2}}{12l_{\Lambda}^{2}}\frac{(X^{-})^{3}}{(P_{+})^{2}},~~~\widetilde{X}^{-}=X^{-}.$ (41)
@@
Thus, we complete the canonical transformation, and the Hamiltonian (35) becomes, in the lapse choice (34), the desired form:
@@
We conclude with three remarks. First, the above symmetry holds for the specific choice of the the lapse function (34). Second, the Hamiltonian constraint is not used to derive the symmetry. Third, the constraint itself appears as one of the generators, $\widetilde{Q}_{+}=H_{g}+\frac{cl_{p}}{2l_{s}^{2}}$. These are involved in understanding the physical meaning of the symmetry (see Secs.VI and VII).
@@
Before closing this section, we here show that the solution of the Hamiltonian (35) is uniquely given by the (A)dS Schwarzschild metric, which provides a physical aspect of this system. In the derivation, we impose the Hamiltonian constraint $H_{g}\approx 0$ as late as possible, to make its role clear.
@@
Combining these and the lapse (34), we can calculate
@@
Here, the formula (53) plays the role of the energy of the system. Indeed, in the case of $\Lambda=0$ and $\epsilon=-1$, it agrees with the Komar mass Wald:1984rg (see Appendix A). $M$ commutes with $H_{g}$ (43) and thus is a gauge-invariant observable. Note that if we introduce another gauge-invariant observable $T\equiv\frac{G\bar{c}^{2}}{\epsilon l_{s}\widetilde{P}_{+}}$, which is associated with $dt_{\infty}\equiv\frac{\bar{c}}{\widetilde{P}_{+}}dt$, we can check that
$\{T,M\}=1,$ (57)
holds, [^4] Similar relations appear in Refs.Kastrup:1993br ; Kuchar:1994zk . and it would be interesting to study this system in terms of $(T,M)$ and consider the effect of the Schrödinger symmetry.
@@ IV Maxwell Field
Here, $\mathcal{H}_{gE}$ is the Hamiltonian constraint, and the canonical variables are $(X^{+},X^{-},X^{A};P_{+},P_{-},P_{A})$.
@@
which agrees with the standard definition of electric charge (see Appendix A) and commutes with $H_{gE}$. In the last line, we employed the constraint (from (70)):
@@ V Multiple Massless Scalar Fields
As another case with matter fields, we discuss the mini-superspace with $n$ massless scalar fields, whose solution corresponds to a generalized Janis-Newman-Winicour (JNW) solution Janis:1968zz and its “interior” Chiou:2008eg , and we show that $(2+n)$ D-Schrödinger symmetry appears. An interesting point here is that the symmetry is obtained by using a different coordinate system $Y^{a}$ in the mini-superspace than the previous one $X^{a}$, together with a different choice of gauge $N$.